# High Torque Density PCB Axial Flux Permanent Magnet Motor for Micro Robots

### v1

@@ I Introduction
In recent years, quasi‑direct‑drive (QDD) actuation has spurred rapid progress across legged machines and manipulators, including quadrupeds [1], bipeds [2], lightweight robot arms [3], and high‑performance grippers [4]. By placing a low‑ratio transmission—or no gearbox at all—between a torque‑dense motor and the joint, QDD robots gain high back‑drivability, low reflected inertia, minimal friction and backlash, and therefore superior force control bandwidth and energy efficiency. These advantages, however, come with two key design constraints: 1) High torque at low speed is required because the gear ratio is small. 2) The joint envelope is usually disc‑shaped—large radius but shallow height—demanding a flat, high‑torque motor topology.
Axial‑flux permanent‑magnet (AFPM) machines naturally satisfy both constraints. Their torque scales with the cube of the rotor radius (versus the square for radial‑flux machines), enabling high torque at low rpm, while their thin pancake geometry fits easily inside robot joints and simplifies heat extraction [5].
@@
To overcome these manufacturing limitations, printed circuit board (PCB)-based winding technology has emerged as an attractive alternative due to its scalable production, cost-effectiveness, and compatibility with microfabrication techniques, making them ideal for micro-robotic applications [6]. Yet off‑the‑shelf PCB motors still suffer from low copper fill factors and therefore limited torque density [7]. Addressing this shortcoming is critical, as it directly influences the motor’s achievable torque and efficiency, both vital metrics in high-performance robotic actuators.
@@ II Design
In this paper, we propose, fabricate, and analyze a $14mm\times 19mm$ motor. Table I summarizes the motor’s specifications. We evaluated several PCB layouts, but with a stator core in place, the performance differences were minimal. Therefore, for ease of fabrication, we adopted a standard PCB layout (see Fig 3(a)). Fig 3(b) shows how a 12-layer module is interconnected using HDI technology—horizontal strips represent the copper layers, and black rectangles denote the interlayer via “stacks” that form the multi-layer coil paths. As indicated in Table I, using IC substrates instead of conventional PCB materials significantly increases the copper fill factor. Notably, our design consists of four series-connected 12-layer modules (totaling 48 layers), which is substantially more than the 6 layers typically found in other PCB motors [8].
@@ III Electromagnetic Analysis and Thermal Analysis
The motor presented in this work is specifically engineered for robotic applications that demand high torque at low rotational speeds and frequent operation under stall conditions. The design methodology follows a structured approach grounded in both electromagnetic [9, 10] and thermal analysis [11].
@@
Similar to [12], the design process begins with an evaluation of magnetic flux density to guide the optimization of the yoke geometry, magnetic material selection, magnet configuration, and coil winding parameters. To prevent magnetic saturation in the stator teeth and yoke, the flux density in the core is constrained to remain below the material’s saturation limit. This is expressed as: $B_{\text{core}}=\frac{\Phi}{A_{\text{core}}}\leq B_{\text{sat}}$ where $\Phi$ is the magnetic flux, $A_{\text{core}}$ is the cross-sectional area of the core (teeth or yoke), and $B_{\text{sat}}$ is the saturation flux density of the core material. Figure 4 presents the estimated no-load magnetic flux density from finite element analysis (FEA) performed using Ansys Maxwell [13]. A high-saturation magnetic material (2.4 T) is selected to enhance magnetic flux capacity and prevent saturation under operational loads.
@@ IV Experiment
We compare our motor against a state-of-the-art commercially available miniature motor [15], which serves as a relevant benchmark despite not being a one-to-one comparison. Our motor is specifically designed for robotic actuators that require high torque at low speeds, whereas the Maxon motor operates at relatively higher speeds. Nevertheless, it represents the closest available reference point for contextualizing the design intent and performance positioning of our prototype.

### v2

@@ I Introduction
In recent years, quasi‑direct‑drive (QDD) actuation has spurred rapid progress across legged machines and manipulators, including quadrupeds [bledt2018cheetah], bipeds [saloutos2023design], lightweight robot arms [zhao2023light], and high‑performance grippers [bhatia2019direct]. By placing a low‑ratio transmission—or no gearbox at all—between a torque‑dense motor and the joint, QDD robots gain high back‑drivability, low reflected inertia, minimal friction and backlash, and therefore superior force control bandwidth and energy efficiency. These advantages, however, come with two key design constraints: 1) High torque at low speed is required because the gear ratio is small. 2) The joint envelope is usually disc‑shaped—large radius but shallow height—demanding a flat, high‑torque motor topology.
Axial‑flux permanent‑magnet (AFPM) machines naturally satisfy both constraints. Their torque scales with the cube of the rotor radius (versus the square for radial‑flux machines), enabling high torque at low rpm, while their thin pancake geometry fits easily inside robot joints and simplifies heat extraction [seo2010design].
@@
To overcome these manufacturing limitations, printed circuit board (PCB)-based winding technology has emerged as an attractive alternative due to its scalable production, cost-effectiveness, and compatibility with microfabrication techniques, making them ideal for micro-robotic applications [wu2012design]. Yet off‑the‑shelf PCB motors still suffer from low copper fill factors and therefore limited torque density [10475522]. Addressing this shortcoming is critical, as it directly influences the motor’s achievable torque and efficiency, both vital metrics in high-performance robotic actuators.
@@ II Design
In this paper, we propose, fabricate, and analyze a $14mm\times 19mm$ motor. Table I summarizes the motor’s specifications. We evaluated several PCB layouts, but with a stator core in place, the performance differences were minimal. Therefore, for ease of fabrication, we adopted a standard PCB layout (see Fig 3(a)). Fig 3(b) shows how a 12-layer module is interconnected using HDI technology—horizontal strips represent the copper layers, and black rectangles denote the interlayer via “stacks” that form the multi-layer coil paths. As indicated in Table I, using IC substrates instead of conventional PCB materials significantly increases the copper fill factor. Notably, our design consists of four series-connected 12-layer modules (totaling 48 layers), which is substantially more than the 6 layers typically found in other PCB motors [wang2021design].
@@ III Electromagnetic Analysis and Thermal Analysis
The motor presented in this work is specifically engineered for robotic applications that demand high torque at low rotational speeds and frequent operation under stall conditions. The design methodology follows a structured approach grounded in both electromagnetic [zhu2000finite, zhou2002field] and thermal analysis [roy2024heat].
@@
Similar to [quan2025design], the design process begins with an evaluation of magnetic flux density to guide the optimization of the yoke geometry, magnetic material selection, magnet configuration, and coil winding parameters. To prevent magnetic saturation in the stator teeth and yoke, the flux density in the core is constrained to remain below the material’s saturation limit. This is expressed as: $B_{\text{core}}=\frac{\Phi}{A_{\text{core}}}\leq B_{\text{sat}}$ where $\Phi$ is the magnetic flux, $A_{\text{core}}$ is the cross-sectional area of the core (teeth or yoke), and $B_{\text{sat}}$ is the saturation flux density of the core material. Figure 4 presents the estimated no-load magnetic flux density from finite element analysis (FEA) performed using Ansys Maxwell [ANSYS_Maxwell_2025R1]. A high-saturation magnetic material (2.4 T) is selected to enhance magnetic flux capacity and prevent saturation under operational loads.
@@ IV Experiment
We compare our motor against a state-of-the-art commercially available miniature motor [maxon], which serves as a relevant benchmark despite not being a one-to-one comparison. Our motor is specifically designed for robotic actuators that require high torque at low speeds, whereas the Maxon motor operates at relatively higher speeds. Nevertheless, it represents the closest available reference point for contextualizing the design intent and performance positioning of our prototype.